OpenStudy (anonymous):

If z varies directly with q and inversely with m, and z=2 when q=3 and m-4, find q when z=5 and m=7

OpenStudy (saifoo.khan):

\[z = \frac{qk}{m}\] Find k. \[2 = \frac{3(k)}{4}\] \[k = \frac83\]

OpenStudy (saifoo.khan):

\[z = \frac83 \times \frac qm\] \[5 = \frac83 \times \frac q7\]Solve for q and get it's value.

OpenStudy (anonymous):

q=13.13